The equation of the straight line with gradient 2 and passing through the point with coordinate (-2, 1) 


1
Expert's answer
2021-08-02T16:47:05-0400

"\text{Gradient = }\frac{y-y_1}{x-x_1} \implies \text{gradient = 2, $y_1 = 1 $ and $x_1 = -2$} \\
\implies \frac{y-1}{x-(-2)} = \frac{y-1}{x+2} = 2 \implies y-1 = 2(x+2) = 2x + 2 \\
\implies y- 1 = 2x + 2 \implies y = 2x + 2+1 = 2x+ 3\\
\text{The required equation of line is: } y = 2x + 3"